# Cross-modal Prompting for Balanced Incomplete Multi-modal Emotion Recognition

### v1

@@ Introduction
However, the aforementioned traditional MER methods are generally based on an assumption that instances from all modalities are fully observed, which is often broken by limitations in data collection, transmission, and storage in real-world scenarios (Gcnet; he2025dual). For instance, background noise might disable the audio modality, and severe accents may lead to failures in speech recognition when constructing the text modality. The incomplete data problem impairs the cross-modal consistency, making it more challenging to achieve accurate emotion recognition.
@@
The recognition accuracies of all methods decline as the MR increases due to the uncertainty and noise introduced by incompleteness. DiCMoR and IDMer restore the original shallow features with generative models, so that traditional MER methods built on complete datasets could be transferred. However, such reconstruction could be noisy and less representative, resulting in relatively inferior performance even at a low MR, with a rapid degradation as MR increases. GCNet employs an early-fusion strategy by concatenating modality features and passing them through an LSTM, which fails to comprehensively investigate the cross-modality correlation at the utterance level and leads to less accurate recognition. MoMKE performs cross-modal communication at the model level by passing features to experts pre-trained independently on single modalities. However, such coarse-grained message passing fails to comprehensively exchange feature-level messaging, and its testing results drop severely when the MR becomes large.
Compared to the existing methods, our delicately designed knowledge propagation module with prompt learning enables more thorough communication across views by mutual information enhancement, which alleviates the side-effect caused by missing instances and improves the overall performance through single-modal performance boosting. Our ComP almost always outperforms other methods on $4$ datasets with different missing rates. Additionally, the increasing number of missing instances has a relatively small impact on our methods than on others, which again validates that our coherency enhancement strategy could effectively extract the task-relevant information even when some instances are missing.

### v2

@@ Introduction
However, the aforementioned traditional MER methods are generally based on an assumption that instances from all modalities are fully observed, which is often broken by limitations in data collection, transmission, and storage in real-world scenarios (Gcnet; he2025dual; zhang2023tensorized). For instance, background noise might disable the audio modality, and severe accents may lead to failures in speech recognition when constructing the text modality. The incomplete data problem impairs the cross-modal consistency, making it more challenging to achieve accurate emotion recognition.
@@
The recognition accuracies of all methods decline as the MR increases due to the uncertainty and noise introduced by incompleteness. DiCMoR and IMDer restore the original shallow features with generative models, so that traditional MER methods built on complete datasets could be transferred. However, such reconstruction could be noisy and less representative, resulting in relatively inferior performance even at a low MR, with a rapid degradation as MR increases. GCNet employs an early-fusion strategy by concatenating modality features and passing them through an LSTM, which fails to comprehensively investigate the cross-modality correlation at the utterance level and leads to less accurate recognition. MoMKE performs cross-modal communication at the model level by passing features to experts pre-trained independently on single modalities. However, such coarse-grained message passing fails to comprehensively exchange feature-level messaging, and its testing results drop severely when the MR becomes large.
Compared to the existing methods, our knowledge propagation module with prompt learning enables more thorough communication across views by mutual information enhancement, which alleviates the side-effect caused by missing instances and improves the overall performance through single-modal boosting. Our ComP almost always outperforms other methods on $4$ datasets with different missing rates. Additionally, the increasing number of missing instances has a relatively small impact on our methods than on others, which again validates that our coherency enhancement strategy could effectively extract the task-relevant information even when some instances are missing.